For how many three-element sets of positive integers is it true that their product is 2013?
Answer by richard1234(7193)

2013 = 3*11*61

If none of the elements is 1, there is one set ({3,11,61})

If exactly one of the elements is 1, there are three sets ({3, 11*61}, {11, 3*61}, {61, 3*11}).

If exactly two of the elements are 1, there is one set ({1, 1, 2013}).

# sets = 1+3+1 = 5.